# Learning to Generalize without Bias for Open-Vocabulary Action Recognition

### v1

@@ Abstract
Leveraging the effective visual-text alignment and static generalizability from CLIP, recent video learners adopt CLIP initialization with further regularization or recombination for generalization in open-vocabulary action recognition in-context. However, due to the static bias of CLIP, such video learners tend to overfit on shortcut static features, thereby compromising their generalizability, especially to novel out-of-context actions. To address this issue, we introduce Open-MeDe, a novel Meta-optimization framework with static Debiasing for Open-vocabulary action recognition. From a fresh perspective of generalization, Open-MeDe adopts a meta-learning approach to improve “known-to-open generalizing” and “image-to-video debiasing” in a cost-effective manner. Specifically, Open-MeDe introduces a cross-batch meta-optimization scheme that explicitly encourages video learners to quickly generalize to arbitrary subsequent data via virtual evaluation, steering a smoother optimization landscape. In effect, the free of CLIP regularization during optimization implicitly mitigates the inherent static bias of the video meta-learner. We further apply self-ensemble over the optimization trajectory to obtain generic optimal parameters that can achieve robust generalization to both in-context and out-of-context novel data. Extensive evaluations show that Open-MeDe not only surpasses state-of-the-art regularization methods tailored for in-context open-vocabulary action recognition but also substantially excels in out-of-context scenarios.
@@ 1 Introduction
Figure 1: Performance comparison (Top-1 Acc (%)) under various open-vocabulary evaluation settings where the video learners except for CLIP are tuned on Kinetics-400 [28] with frozen text encoders. The satisfying in-context generalizability on UCF101 [44] (a) can be severely affected by static bias when evaluating on out-of-context SCUBA-UCF101 [31] (b) by replacing the video background with other images.
Open-vocabulary Action Recognition (OVAR) aims to identify test videos whose classes are not previously encountered during the training phase, which challenges the generalization and zero-shot capabilities of the video learners [59, 3, 53]. Recently, the emergence of image-based visual-language (I-VL) pre-training, such as CLIP [40] and ALIGN [25], has shown promising zero-shot inference in image-based tasks. Inspired by this success, recent attempts [3, 50, 7, 35, 37] have been made to adapt CLIP for general action recognition via additional temporal modeling following the “pre-train, prompt and fine-tune” paradigm [49]. Broadly, these video learners optimize the learnable parameters from the start point of CLIP, pursuing decent performance on the training videos, known as standard fine-tuning objectives. However, adapting CLIP to the video domain, especially for OVAR, is extremely challenging, as the video learners with standard fine-tuning objectives often lead to overfitting, which achieves improved specialization at the cost of generalization degradation.
@@
Based on this insight, we propose Open-MeDe, a novel Meta-learning based framework with static Debiasing for in-context and out-of-context Open-vocabulary action recognition. Meta-learning, also known as “learning to learn”, incorporates virtual evaluation during the training process for better generalization [19, 36, 1]. In our meta-learning scheme, the “learning to generalize” process is enhanced by naturally treating sequences of adjacent batches sampled from the training set as a distribution of tasks. More concretely, our procedure optimizes the video learner to obtain fast weights by gradient descent updates on the current batch (i.e., meta training), while evaluating the subsequent batch (i.e., meta testing) based on fast weights of the learner, which mimics a known-to-open task. Based on the evaluation performance in meta testing, our procedure can further optimize the learner to obtain more generalizable video-specific knowledge against inherent known and static biases. In effect, this cross-batch meta-optimization formulates a meta-learner free of CLIP regularization, thereby facilitating smoother optimization and robust video representation learning for fast known-to-open generalizing, thus enhancing image-to-video debiasing. Tailored to the optimization trajectory of the video learner, we further employ self-ensemble stabilization, i.e., Gaussian Weight Average (GWA), to derive generic optima for robust generalization at open-vocabulary test time. Overall, while integrating the same video learner, our model-agnostic Open-MeDe outperforms existing regularization-based methods, which strikes a promising balance on in-context and out-of-context generalization settings (Fig. 1).
@@ 2.2 Meta-learning
Rather than directly learning from experiences, with the goal of learning to learn, meta-learning can quickly generalize to new tasks by leveraging prior learning abilities [21]. As the representative works in meta-learning, MAML [19] boasts simplicity and has actively driven the development of the gradient-based methods in few-shot learning. Recently, meta-learning techniques have also been explored in zero-shot learning [46, 38, 33, 22], which typically perform episode-wise training by dividing the training set into support and query sets with different classes distributions. Targeting long-tailed issues within closed-set video scene generation, MVSGG [55] employs meta-learning across various types of tasks w.r.t. certain conditional biases through meticulous structuring of training data. However, these approaches are often prone to meta-overfitting due to insufficient meta tasks and limited application scopes of generalization. Differently, our proposed meta-optimization scheme naturally mimics diverse known-to-open tasks incurring no additional computational overhead. This tackles ubiquitous challenges in video understanding beyond closed-set and in-context settings, i.e., mitigating static bias of video learners for open-vocabulary generalization.
@@ 3.1 Preliminaries
Action recognition with CLIP-based video learner. Pre-training on large-scale image-text data based on contrastive learning, CLIP learns separate uni-modal encoders for image and text, embedding them into a joint feature space, respectively. Consider a CLIP-based video learner with a ViT architecture [17], that incorporates temporal modeling for video understanding [53, 49, 51, 59, 56, 50]. Next, we present the standard vision-only fine-tuning paradigm that applies such a video learner $f_{\theta_{v}}$ with a frozen text encoder $f_{\theta_{t}}$ to action recognition. Specifically, given a video clip $V_{i}$, and a candidate action label $T_{j}\in\mathcal{Z}_{tr}$ described in predefined textual templates (e.g., “a video of {action}”) from the training set $\mathcal{D}_{tr}$, the similarity is calculated as:
@@ 3.2.1 Cross-batch meta-optimization
Meta testing. After meta training on the support batch, we then scheme a virtual testing process, leveraging the query batch $\mathcal{Q}=\{V_{i},T_{i}\}^{B}$, where $\mathcal{S}\cap\mathcal{Q}=\emptyset$, to evaluate the generalization performance of the base learner $f_{\theta_{i}^{\prime}}$. Formally, we measure the known-to-open performance on $\mathcal{T}_{i}$ by calculating the class-specific loss concerning the query data as:
@@
Here, the formulation closely relates to the standard fine-tuning process, which aims to obtain decent class-specific performance for all training batches. Differently, this step merely evaluates the intermediary base learner for its known-to-open generalizability on each task, due to the original parameters $\theta$ remaining immune to the task-specific updates. Hence, it can be used to provide feedback on what video-specific knowledge should be learned in the sense that the learner can derive the robust generalization across different class distributions against inherent known and static biases in the following meta-optimization.
@@
where $N$ is the batch size of the task for meta-optimization. Since the MAML meta-gradient update needs to differentiate through the optimization process (i.e., a gradient by a gradient), it’s not an ideal solution where we need to optimize a large number of tasks during the training phase. Therefore, we opt for the one-step update approximation by dropping the backward pass of $\theta\leftarrow\theta^{\prime}_{i}$, where Eq. (9) can be rewritten as:
@@
Frozen CLIP [40] 62.3 53.4 57.5 53.3 46.8 49.8 78.5 63.6 70.3 4.9 5.3 5.1
ActionCLIP [49] 61.0 46.2 52.6 69.1 37.3 48.5 90.1 58.1 70.7 13.3 10.1 11.5
X-CLIP [35] 74.1 56.4 64.0 69.4 45.5 55.0 89.9 58.9 71.2 8.5 6.6 7.4
VPT [26] 69.7 37.6 48.8 46.2 16.0 23.8 90.5 40.4 55.8 8.3 5.3 6.4
ST-Adapter [37] 74.6 62.0 67.3 65.3 48.9 55.9 85.5 76.8 80.9 9.3 8.4 8.8
ViFi-CLIP [41] 76.4 61.1 67.9 73.8 53.3 61.9 92.9 67.7 78.3 16.2 12.1 13.9
Open-VCLIP $\ast$ [51] 76.3 62.3 68.6 70.2 50.2 58.5 94.6 77.2 85.0 15.9 10.8 12.9
FROSTER $\dagger$ [23] 76.0 61.9 68.3 70.0 49.9 58.3 94.3 76.9 84.7 15.5 10.3 12.4
@@ 3.2.2 Gaussian self-ensemble stabilization
Typically, training the video learner for longer iterations to gain specialization on the supervised tasks comes with the risk of diminished plasticity and generalizability. Model patching [24, 52, 43, 51] of weight ensembling has been shown to improve both the performance and generalization. Given that the fine-tuning videos are limited in class-specific knowledge, while the open-vocabulary tasks are unconstrained, the static generalizable flexibility derived from large-scale I-VL pre-training should be scrupulously exploited to enhance the adaptation of the video learner while minimizing the impact of static bias. Therefore, we further incorporate self-ensemble stabilization tailored to the video learner over its optimization trajectory, which utilizes the knowledge from previous training iterations for a generalizable solution. In a fine-tuning procedure of $R$ epochs with $l$ step length for each, the learner’s optimization trajectory is represented by $\{\theta_{t}\}_{t=0}^{R}$, where $\theta_{0}$ is the pre-trained weights. The self-ensemble averages the weights of the learner as:
$\theta_{\texttt{WA}}=\sum_{t=0}^{R}\frac{w_{t}}{\sum_{i=0}^{R}w_{i}}\cdot\theta_{t},$ (11)
where $w_{t}$ specifies the weight contributed by the parameters at $t$-th epoch. Intuitively, during the early fine-tuning epochs (i.e., at a smaller epoch $t$), the video learner lacks the maturity to effectively capture video-specific knowledge while still retaining substantial static-related orientation from large-scale pre-training, which introduces vulnerable information for temporal understanding. Conversely, the parameters at the last few epochs (i.e., at a larger epoch $t$) have integrated more video-specific knowledge, highly featuring the supervised downstream task distribution, whereas the plasticity of the unconstrained zero-shot capability is not guaranteed. As both sides degrade the final open-vocabulary generalizability, we aim to weaken the contribution of the parameters near the initial and terminal epochs by employing a distribution prior, resulting in a generic optima for robust generalization.
@@ 13             $t\leftarrow t+1;\theta_{t}\leftarrow\theta$
Here, we exclude the integration of CLIP weights $\theta_{0}$ for the purpose of static debiasing. $\mu$ and $\sigma^{2}$ are hyper-parameters for the distribution, and in practice, we determine the value of $\mu$ according to the epoch number. Then, we perform normalization towards the weights of total epochs to achieve $\sum_{t=1}^{R}w_{t}=1$. We also formulate GWA as a moving average to avoid increasing the storage cost of saving multiple snapshots of the parameters by updating the average of current learner $\theta_{t}$ on the fly (i.e., at epoch $t$) as:
@@ 3.3 Algorithm overview
We present the overall training procedure of the proposed model-agnostic Open-MeDe in Algorithm 1. The video learner is fine-tuned on training videos based on our cross-batch meta-optimization scheme cost-effectively. And the Gaussian self-ensemble stabilization is performed on the video learner via our GWA for robust generalization under open-vocabulary settings.
@@ 4.1 Experimental Setup
Implementation details. Generally, we use the official CLIP ViT-B/16 backbone for all experiments, and our video learner is the adaptation of the CLIP model follows [51], unless stated otherwise. During our meta-optimization process, we construct a batch of 4 tasks, each task contains 8 support and query samples from the training set. The learning rates of inner and outer loops for support batches i.e., $\alpha$, and $\beta$, are synchronized with the initial value of $3.33\times 10^{-6}$ and decay to $3.33^{-8}$ utilizing the AdamW [34] optimizer following a cosine decay scheduler, while the hyperparameter $\delta$ for query batches is set to $1.67\times 10^{-3}$. For cross-dataset evaluation, we warm up the training on the K400 dataset for the first 2 epochs and further fine-tune the video learner for 20 epochs. For base-to-novel evaluation, we train the learner for 12 epochs with the first two warm-up epochs on training data. During inference, we use 3 temporal and 1 spatial views per video and linearly aggregate the recognition results. See Appendix for more experimental details about evaluation protocols and our implementations.
@@ Acknowledgments
This work is supported by National Natural Science Foundation of China (No. 62376217, 62301434), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), Key R&D Project of Shaanxi Province (No. 2023-YBGY-240), and Young Talent Fund of Association for Science and Technology in Shaanxi, China (No. 20220117).
@@ B.1 Additional Evaluations and Ablation Studies
Table 9: Effect of the randomness of the batch sampler for cross-batch meta-optimization. We evaluate both in-context and out-of-context recognition (marked with $\star$) performances. HM: harmonic mean.
@@
Plain 77.7 28.2 41.4
78.5 ($+0.8\%$) 28.3 ($+0.1\%$) 41.6 ($+0.2\%$)
Meta Cross-batch 82.5 28.9 42.8
83.7 ($+1.2\%$) 33.5 ($+4.6\%$) 47.8($+5.0\%$)
Effect of randomness of the batch sampler for cross-batch meta-optimization. To verify the efficacy of constructing tasks across batches with different inherent label distributions, we further conduct several additional studies about the sampling randomness during cross-batch meta-optimization. As shown in Tab. 9, the randomness of the batch sampler is indeed an important factor to bring out the best of our cross-batch meta leaner, which improves the overall generalization greatly ($+5.0\%$ of harmonic mean) especially for out-of-context performance ($+4.6\%$ on UCF-SCUBA). However, Plain learner shows insensibility to the sampling randomness, experiencing negligible growth of generalization performance. Without shuffling the batch sampler, our method still outperforms the non-shuffle plain learner by $+1.4\%$ of HM results.

### v2

@@ Abstract
Leveraging the effective visual-text alignment and static generalizability from CLIP, recent video learners adopt CLIP initialization with further regularization or recombination for generalization in open-vocabulary action recognition in-context. However, due to the static bias of CLIP, such video learners tend to overfit on shortcut static features, thereby compromising their generalizability, especially to novel out-of-context actions. To address this issue, we introduce Open-MeDe, a novel Meta-optimization framework with static Debiasing for Open-vocabulary action recognition. From a fresh perspective of generalization, Open-MeDe adopts a meta-learning approach to improve “known-to-open generalizing” and “image-to-video debiasing” in a cost-effective manner. Specifically, Open-MeDe introduces a cross-batch meta-optimization scheme that explicitly encourages video learners to quickly generalize to arbitrary subsequent data via virtual evaluation, steering a smoother optimization landscape. In effect, the free of CLIP regularization during optimization implicitly mitigates the inherent static bias of the video meta-learner. We further apply self-ensemble over the optimization trajectory to obtain generic optimal parameters that can achieve robust generalization to both in-context and out-of-context novel data. Extensive evaluations show that Open-MeDe not only surpasses state-of-the-art regularization methods tailored for in-context open-vocabulary action recognition but also substantially excels in out-of-context scenarios. Code is released at https://github.com/Mia-YatingYu/Open-MeDe.
@@ 1 Introduction
Figure 1: Performance comparison (Top-1 Acc (%)) under various open-vocabulary evaluation settings where the video learners except for CLIP are tuned on Kinetics-400 [29] with frozen text encoders. The satisfying in-context generalizability on UCF101 [46] (a) can be severely affected by static bias when evaluating on out-of-context UCF-SCUBA [33] (b) by replacing the video background with other images.
Open-vocabulary action recognition (OVAR) aims to identify test videos whose classes are not previously encountered during the training phase, which challenges the generalization and zero-shot capabilities of the video learners [62, 3, 56, 52]. Recently, the emergence of image-based visual-language (I-VL) pre-training, such as CLIP [42] and ALIGN [26], has shown promising zero-shot inference in image-based tasks. Inspired by this success, recent attempts [3, 53, 8, 37, 39, 4] have been made to adapt CLIP for general action recognition via additional temporal modeling following the “pre-train, prompt and fine-tune” paradigm [51]. Broadly, these video learners optimize the learnable parameters from the start point of CLIP, pursuing decent performance on the training videos, known as standard fine-tuning objectives. However, adapting CLIP to the video domain, especially for OVAR, is extremely challenging, as the video learners with standard fine-tuning objectives often lead to overfitting, which achieves improved specialization at the cost of generalization degradation.
@@
Based on this insight, we propose Open-MeDe, the first Meta-learning based framework with static Debiasing for in-context and out-of-context Open-vocabulary action recognition. Meta-learning, also known as “learning to learn”, incorporates virtual evaluation during the training process for better generalization [20, 38, 1]. In our meta-learning scheme, the “learning to generalize” process is enhanced by naturally treating sequences of adjacent batches sampled from the training set as a distribution of tasks. More concretely, our procedure optimizes the video learner to obtain fast weights by gradient descent updates on the current batch (i.e., meta training), while evaluating the subsequent batch (i.e., meta testing) based on fast weights of the learner, which mimics a known-to-open task. Based on the evaluation performance in meta testing, our procedure can further optimize the learner to obtain more generalizable video-specific knowledge against inherent known and static biases. In effect, this cross-batch meta-optimization formulates a meta-learner free of CLIP regularization, thereby facilitating smoother optimization and robust video representation learning for fast known-to-open generalizing, thus enhancing image-to-video debiasing. Tailored to the optimization trajectory of the video learner, we further employ self-ensemble stabilization, i.e., Gaussian Weight Average (GWA), to derive generic optima for robust generalization at open-vocabulary test time. Overall, while integrating the same video learner, our model-agnostic Open-MeDe outperforms existing regularization-based methods, which strikes a promising balance on in-context and out-of-context generalization settings (Fig. 1).
@@ 2.2 Meta-learning
Rather than directly learning from experiences, with the goal of learning to learn, meta-learning can quickly generalize to new tasks by leveraging prior learning abilities [22]. As the representative works in meta-learning, MAML [20] boasts simplicity and has actively driven the development of the gradient-based methods in few-shot learning. Recently, meta-learning techniques have also been explored in zero-shot learning [48, 40, 35, 23] and domain adaptation [32], which typically perform episode-wise training by dividing the training set into support and query sets with different classes distributions. Targeting long-tailed issues within closed-set video scene generation, MVSGG [58] employs meta-learning across several manually predefined task types, which are partitioned based on specific conditional biases in the training data. However, these approaches are often prone to meta-overfitting due to insufficient meta tasks and limited application scopes of generalization. Differently, our work tackles ubiquitous challenges in video understanding beyond closed-set and in-context settings, i.e., mitigating static bias of video learners for open-vocabulary generalization. To the best of our knowledge, we are the first to directly integrate the mini-batch training mechanism with meta-learning to naturally mimic diverse known-to-open tasks utilizing cross-batch data without additional computational overhead.
@@ 3.1 Preliminaries
Action recognition with CLIP-based video learner. Consider a CLIP-based video learner with a ViT architecture [18], that incorporates temporal modeling for video understanding [56, 51, 54, 62, 59, 53]. Next, we present the standard vision-only fine-tuning paradigm that applies such a video learner $f_{\theta_{v}}$ with a frozen text encoder $f_{\theta_{t}}$ to action recognition. Specifically, given a video clip $V_{i}$, and a candidate action label $T_{j}\in\mathcal{Z}_{tr}$ described in predefined textual templates (e.g., “a video of {action}”) from the training set $\mathcal{D}_{tr}$, the similarity is calculated as:
@@ 3.2.1 Cross-batch meta-optimization
Meta testing. After meta training on the support batch, we then scheme a virtual testing process, leveraging the query batch $\mathcal{Q}=\{V_{i},T_{i}\}^{B}$, where $\mathcal{S}\cap\mathcal{Q}=\emptyset$, to evaluate the generalization performance of the base learner $f_{\theta_{i}^{\prime}}$. Formally, we measure the known-to-open performance on $\mathcal{T}_{i}$ by calculating the class-specific loss w.r.t. the query data as:
@@
Here, the formulation closely relates to the standard fine-tuning process, which aims to obtain decent class-specific performance for all training batches. Differently, this step merely evaluates the intermediary base learner for its known-to-open generalizability on each task, due to the original parameters $\theta$ remaining immune to the task-specific updates. Hence, it can be used to provide feedback for the learner on what video-specific knowledge should be learned to derive the robust generalization against inherent known and static biases in the following meta-optimization.
@@
where $N$ is the batch size of the task for meta-optimization. Since the MAML meta-gradient update needs to differentiate through the optimization process (i.e., a gradient by a gradient), it’s not an ideal solution where we need to optimize a large number of tasks during the training phase. Therefore, we opt for the one-step update approximation by dropping the backward pass of $\theta\leftarrow\theta^{\prime}_{i}$ as:
@@
Frozen CLIP [42] ICML’21 62.3 53.4 57.5 53.3 46.8 49.8 78.5 63.6 70.3 4.9 5.3 5.1
ActionCLIP [51] arXiv’21 61.0 46.2 52.6 69.1 37.3 48.5 90.1 58.1 70.7 13.3 10.1 11.5
X-CLIP [37] ECCV’22 74.1 56.4 64.0 69.4 45.5 55.0 89.9 58.9 71.2 8.5 6.6 7.4
VPT [27] ECCV’22 69.7 37.6 48.8 46.2 16.0 23.8 90.5 40.4 55.8 8.3 5.3 6.4
ST-Adapter [39] NeurIPS’22 74.6 62.0 67.3 65.3 48.9 55.9 85.5 76.8 80.9 9.3 8.4 8.8
ViFi-CLIP [43] CVPR’23 76.4 61.1 67.9 73.8 53.3 61.9 92.9 67.7 78.3 16.2 12.1 13.9
Open-VCLIP $\ast$ [54] ICML’23 76.3 62.3 68.6 70.2 50.2 58.5 94.6 77.2 85.0 15.9 10.8 12.9
FROSTER $\dagger$ [24] ICLR’24 76.0 61.9 68.3 70.0 49.9 58.3 94.3 76.9 84.7 15.5 10.3 12.4
@@ 3.2.2 Gaussian self-ensemble stabilization
Typically, training the video learner for longer iterations to gain specialization on the supervised tasks comes with the risk of diminished plasticity and generalizability. Model patching [25, 55, 45, 54] of weight ensembling has been shown to improve both the performance and generalization. Given that the fine-tuning videos are limited in class-specific knowledge, while the open-vocabulary tasks are unconstrained, the static generalizable flexibility derived from large-scale I-VL pre-training should be scrupulously exploited to enhance the adaptation of the video learner while minimizing the impact of static bias. Therefore, we further incorporate self-ensemble stabilization tailored to the video learner over its optimization trajectory, which utilizes the knowledge from previous training iterations for a generalizable solution. In a fine-tuning procedure of $R$ epochs with $l$ step length for each, the learner’s optimization trajectory is represented by $\{\theta_{t}\}_{t=1}^{R}$, and $\theta_{0}$ is the pre-trained weights. The self-ensemble averages the weights of the learner as:
$\theta_{\texttt{WA}}=(1-\sum_{t=1}^{R}\alpha_{t})\cdot\theta_{0}+\sum_{t=1}^{R}{\alpha_{t}}\cdot\theta_{t},$ (11)
where $\alpha_{t}\in[0,1]$ specifies the weights contributed by the parameters at $t$-th epoch. Intuitively, during the early fine-tuning epochs (i.e., at a smaller epoch $t$), the video learner lacks the maturity to effectively capture video-specific knowledge while still retaining substantial static-related orientation from large-scale pre-training, which introduces vulnerable information for temporal understanding. Conversely, the parameters at the last few epochs (i.e., at a larger epoch $t$) have integrated more video-specific knowledge, highly featuring the supervised downstream task distribution, whereas the plasticity of the unconstrained zero-shot capability is not guaranteed. As both sides degrade the final open-vocabulary generalizability, we aim to weaken the contribution of the parameters near the initial and terminal epochs by employing a distribution prior, resulting in a generic optima for robust generalization.
@@ 13             $t\leftarrow t+1;\theta_{t}\leftarrow\theta$
Here, we exclude the integration of CLIP weights $\theta_{0}$ for the purpose of static debiasing. $\mu$ and $\sigma^{2}$ are hyper-parameters for the distribution, and in practice, we determine the value of $\mu$ according to the epoch number. Then, we perform normalization towards the weights of total epochs i.e., $\alpha_{t}=\frac{w_{t}}{\sum_{i=1}^{R}w_{i}}$. We also formulate GWA as a moving average to avoid increasing the storage cost of saving multiple snapshots of the parameters by updating the average of current learner $\theta_{t}$ on the fly (i.e., at epoch $t$) as:
@@ 3.3 Algorithm overview
We present the overall training procedure of the proposed model-agnostic Open-MeDe in Algorithm 1. The video learner is fine-tuned based on training videos based on our cross-batch meta-optimization scheme cost-effectively. And the Gaussian self-ensemble stabilization is performed on the video learner via our GWA for robust generalization under open-vocabulary settings.
@@ 4.1 Experimental Setup
Implementation details. Generally, we use the official CLIP ViT-B/16 backbone for all experiments, and our video learner is the adaptation of the CLIP model follows [54], unless stated otherwise. During our meta-optimization process, we construct a batch of 4 tasks, each task contains 8 support and query samples from the training set. The learning rates of inner and outer loops for support batches i.e., $\alpha$, and $\beta$, are synchronized with the initial value of $3.33\times 10^{-6}$ and decay to $3.33^{-8}$ utilizing the AdamW [36] optimizer following a cosine decay scheduler, while the hyperparameter $\delta$ for query batches is set to $1.67\times 10^{-3}$. For cross-dataset evaluation, we warm up the training on the K400 dataset for the first 2 epochs and further fine-tune the video learner for 20 epochs. For base-to-novel evaluation, we train the learner for 12 epochs with the first two warm-up epochs on training data. During inference, we use 3 temporal and 1 spatial views per video and linearly aggregate the recognition results. See Appendix for experimental details.
@@ Acknowledgments
This research is supported by the National Natural Science Foundation of China (No. 62376217, 62273347, and 62301434), and the Young Elite Scientists Sponsorship Program by CAST (No. 2023QNRC001).
@@ B.1 Additional Evaluations and Ablation Studies
Table 7: Effect of the meta-optimization and Gaussian weight average (GWA) components in Open-MeDe. $\Delta$ denotes the performance gains of different schemes over the baseline. Our Open-MeDe is highlighted in gray.
Meta
optimization
GWA UCF $\Delta_{\text{UCF}}$ UCF-SCUBA $\Delta_{\text{UCF-SCUBA}}$
78.5 - 28.3 -
83.2 $+$ 4.7 32.1 $+$ 3.8
82.3 $+$ 3.8 30.7 $+$ 2.4
83.9 $+$ 5.4 33.5 $+$ 5.2
Effect of individual strategies in Open-MeDe. In Tab. 7, we conduct ablation experiments to study effects of the core strategies in Open-MeDe i.e. the cross-batch meta-optimization and GWA stabilization. Using only meta-optimization or GWA yields improvements of $+4.7\%/3.8\%$ and $+3.8\%/2.4\%$ over the plain learner on UCF / UCF-SCUBA, respectively. This indicates that meta-optimization substantially enhances generalization across both open-vocabulary settings in improving model’s robustness, compared to GWA. These two components complement each other effectively, achieving substantial gains of $+5.4\%/5.2\%$. Their integration leads to consistent improvements across diverse scenarios.
@@
Table 10: Effect of the randomness of the batch sampler for cross-batch meta-optimization. The “similar” sampler denotes the usage of the most semantically similar classes across adjacent batches. We evaluate both in-context and out-of-context recognition (marked with $\star$) performances. HM: harmonic mean. Our default settings and results are highlighted in gray.
@@
Plain shuffle 78.5 28.3 41.6
initial 77.7 ($\downarrow$ 0.8) 28.2 ($\downarrow$ 0.1) 41.4 ($\downarrow$ 0.2)
Meta Cross-batch shuffle 83.7 33.5 47.8
initial 82.5 ($\downarrow$ 1.2) 28.9 ($\downarrow$ 4.6) 42.8 ($\downarrow$ 5.0)
similar 82.7 ($\downarrow$ 1.0) 30.9 ($\downarrow$ 2.6) 45.0 ($\downarrow$ 2.8)
Effect of randomness of the batch sampler for cross-batch meta-optimization. To verify the efficacy of constructing tasks across batches with different inherent label distributions, we further conduct several additional studies about the sampling randomness during cross-batch meta-optimization. As shown in Tab. 10, the randomness of the batch sampler is indeed an important factor to bring out the best of our cross-batch meta leaner, which improves the overall generalization greatly ($+5.0\%$ of harmonic mean) especially for out-of-context performance ($+4.6\%$ on UCF-SCUBA). However, plain learner shows insensibility to the sampling randomness, experiencing negligible growth of generalization performance. Without shuffling the batch sampler, our method still outperforms the non-shuffle plain learner by $+1.4\%$ of HM results. By using the most semantically similar classes across support and query batches, it brings a relative performance decline of $0.99\%$ and $2.64\%$ on UCF and UCF-SCUBA, respectively. We speculate that it amplifies inter-task semantic distribution shifts hindering cross-task generalization. In contrast, the default ensures consistent and balanced distributions of both inter- and intra-task variance.
@@
Figure 5: Comparison between Open-MeDe and OpenVCLIP across $22$ classes on UCF-SCUBA.
Analysis of class-wise performance. In Fig. 5, we further present the improvements of our Open-MeDe over OpenVCLIP on out-of-context UCF-SCUBA across 22 novel classes. It can be observed that Open-MeDe wins across $19$ of the $22$ classes. The improved categories involve localized motions, where most of the static content is misinterpreted by irrelevant context noise on UCF-SCUBA. We attribute these gains primarily to its effectiveness in static debiasing and capturing fine-grained dynamics. However, its performance is slightly compromised in cases involving team sports or rapid shifts in spatial locations.